RSA census bureau data indicates that the average price of color TV sets can be expressed as linear function of number of sets sold N(thousands) in addition as N increase by 1000 price dropped by R10.40 and when 6485(thousands) sets were sold the average price sets was R504.39
Write equation of line defined by this information

Answer by Theo(13342)


your problem is asking for the linear equation regarding the relationship between the increase in number of sets sold and the price for each set.

this can be found as follows:

the general form of a linear equation is y = mx + b.

m is the slope and b is the y-intercept.

if we let y = the price for each set, and x = the number of sets, then we can find the slope because the change in y divided by the change in x becomes -10.4 / 1000 = -.0104 which is the slope of the linear equation.

this is the value of m in the general form of the linear equation.

the general equation of y = mx + b becomes y = -.0104 * x + b

next we want to find b, which is the y-intercept.

we can do that by replacing y with 504.39 and x with 6485 to get:

504.39 = -.0104 * 6485 + b

now we can solve for b to get:

b = 504.39 + .0104 * 6485

this makes b = 571.834

this makes the general equation of y = mx + b equal to y = -.0104 * x + 571.834

that's your equation.

the graph of that equation looks like this:

$$$

if you let n = x and p = y, then your equation becomes:

p = -.0104 * n + 571.834

when n = 6485, this equation becomes p = -.0104 * 6485 + 571.834 which becomes p = 504.39